# BSTcontrol

### v2

@@ 2.5 Fine-tuning
Fine-tuning pre-trained models with domain-specific data enhances performance for particular communication environments, which is essential for 6G and beyond as it eliminates the need for retraining from scratch. However, fine-tuning large AI models on edge nodes faces challenges beyond memory usage, such as high computational demands, energy efficiency, and network latency [67]. The frequent exchange of gradients or parameters can also strain distributed learning frameworks. Additionally, techniques are necessary to adapt base models to high-level commands or specialized domains like telecommunications. To overcome these challenges, several promising techniques are introduced.
@@ 2.5.1 Parameter-Efficient Fine-Tuning
Low-rank adaptation (LoRA) [68] adapts large pre-trained models by introducing low-rank updates to the weight matrices, freezing the original model parameters and training only the low-rank matrices, denoted as
@@
where $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ represents a pre-trained weight matrix which remains frozen during adaptation. The update $\Delta\mathbf{W}$ is constrained to be low-rank by decomposing it into the product of two smaller matrices, $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$, where the rank $r\ll\min(d,k)$. Only the parameters of $\mathbf{A}$ and $\mathbf{B}$ are trained, drastically reducing the number of trainable parameters compared to updating the full $\mathbf{W}_{0}$ or an unconstrained $\Delta\mathbf{W}$. This reduces the number of trainable parameters, making it ideal for resource-constrained environments. Variants like sparse LoRA [69] enhance efficiency using sparse expert modules, while weight-decomposed LoRA [70] improves learning capacity by decomposing weights into magnitude and direction. LoRA also accelerates convergence in complex wireless datasets by simplifying the optimization landscape, preventing overfitting, and enabling faster adaptation to dynamic environments, improving model efficiency for tasks like traffic prediction and resource allocation.
@@
Prefix tuning [71] is a parameter-efficient method that adds a continuous, task-specific prefix sequence to the input or hidden layers of a model. This prefix sequence is not made up of real tokens but instead consists of learnable parameters. The prefix can influence the behavior of model without modifying the entire weights. This approach allows models to retain their pre-trained parameters while optimizing only the task-specific prefix, making it memory-efficient and computationally light.
@@
Prompt tuning [72] can be viewed as a simplified version of prefix tuning, where only a learnable prefix is added to the input text. This method has shown remarkable scaling capabilities; for sufficiently large models, prompt tuning alone can achieve performance comparable to full fine-tuning. By optimizing just the prompt parameters, prompt tuning provides an efficient way to adapt pre-trained models to specific tasks, minimizing the number of parameters that need to be updated. In wireless networks, prompt tuning can be used to efficiently optimize models for various real-time applications, reducing memory and computational costs, and ensuring practical deployment on edge devices.
@@ 2.5.3 Instruction Tuning
Instruction tuning [73] adapts a pre-trained model to better understand and follow natural language instructions across a wide range of tasks. Rather than training on task-specific data alone, the model learns from diverse instruction-response pairs, improving generalization and flexibility. In wireless networks, this allows operators to control or query the model using high-level commands such as “allocate bandwidth efficiently” or “detect congestion hotspots.” This approach reduces the need for hardcoded logic and enables more human-aligned behavior in communication environments. Instruction tuning is especially useful when tasks vary frequently or lack large labeled datasets.
@@ 2.6.1 Prompt-based Inference
LLMs can solve tasks without parameter updates by using prompts with instructions and examples, as in-context learning [37]. Instead of fine-tuning, the model generalizes from a few input-output demonstrations, enabling zero-shot and few-shot inference. This allows a single frozen model to adapt to various tasks by changing the prompt. Prompt engineering guides the model by designing text prompts, which can include system instructions, user queries, or domain-specific examples. This helps the model understand wireless contexts. Combined with instruction tuning [73], prompt engineering enables large models to support diverse use cases such as configuration, protocol analysis, and troubleshooting.
2.6.2 Chain-of-Thought
While basic prompts can handle simple queries, complex wireless problems benefit from multi-step reasoning. Chain-of-thought (CoT) prompting encourages the model to reason through intermediate steps before producing a final a answer [74]. By appending phrases like reason step by step to the prompt, CoT enables more accurate and explainable results on tasks that require logical or numerical inference. In wireless systems, CoT is useful for anomaly detection, interference management, and resource optimization by guiding the model to iteratively explore the problem.
@@ 2.6.3 Task Planning
Task planning combines prompt engineering and CoT to enable structured workflows. Instead of answering a single query, the model decomposes a high-level goal into sub-tasks, solves each step, and compiles the result. For example, optimizing a network for a large event may involve collecting performance metrics, identifying bottlenecks, and proposing reconfiguration actions. Frameworks like HuggingGPT [75] demonstrate how LLMs can orchestrate such workflows by breaking down tasks and invoking external tools. This pipeline-style planning is well-suited to autonomous wireless networks that involve sensing, analysis, and control.
@@ 2.6.4 Retrieval Augmented Generation
Retrieval augmented generation (RAG) is an approach that enhances inference by coupling it with an external knowledge retriever [76]. In a RAG pipeline, a retriever first finds relevant documents or facts from a domain-specific datastore based on the query. Those retrieved text snippets are then provided as additional context to the generator, which integrates this evidence into its response. This mechanism effectively grounds the output in authoritative data, improving factual accuracy and reducing hallucinations [76]. For example, TelecomRAG [77] applies this paradigm to telecom standards by building a knowledge base from 3GPP specification documents and enabling an LLM to answer protocol questions with precise, verifiable references. In wireless communications, RAG is useful for tasks like dynamic protocol lookup, fault diagnosis, and network troubleshooting [78]. An LLM agent can query a repository of network logs, configuration databases, or radiofrequency (RF) sensor readings and reason about network states in real-time. This yields grounded and trustworthy answers, especially important in rapidly evolving domains like wireless communications.
2.6.5 Mixture-of-Models
@@
Mixture-of-models uses multiple specialized models to tackle complex tasks by combining their expertise [79]. In practice, queries are distributed to a set of models, each offering insights based on their specialized knowledge. The outputs are then evaluated and synthesized by a coordinating model, which selects or combines the most accurate responses. This approach helps improve accuracy, mitigate biases, and ensure more reliable decision-making in domain-specific applications, such as telecommunications [80], where precise expertise is critical.
@@ 2.7 Multimodal Models and Multimodal Alignment
The increasing complexity of wireless systems with their need to handle various types of data modalities (e.g., signals, images, text, and audio) has driven the development of multimodal models. These models integrate diverse modalities into a unified system, enabling more holistic reasoning and decision-making. By processing and aligning information across multiple types of input, multimodal models unlock powerful new applications in wireless communication, such as cross-modal inference, environment-aware communication, and sensor fusion. This section explores the characteristics and applications of multimodal models, as well as the importance of multimodal alignment in enabling seamless integration and understanding of diverse data types.
2.7.1 Multimodal Models
Multimodal models integrate information from multiple data types, such as text, images, audio, and RF signals. Architectures like contrastive language-image pre-training [81], DALL·E [82] and stable diffusion use modality-specific encoders followed by cross-modal fusion layers to process joint representations. Models such as Hunyuan-DiT [49] use transformer-based diffusion for image synthesis, while Qwen-VL [83] processes images and text for tasks like captioning and visual question answering. In wireless communication, multimodal models enable applications like vision-assisted beamforming, sensor fusion, and cross-modal semantic inference. These models allow AI systems to reason over visual, textual, and RF inputs in unified ways, supporting tasks like signal classification or environment-aware communication.
2.7.2 Multimodal Alignment
Multimodal alignment ensures that different modalities are mapped into a shared semantic space. This allows the system to associate related concepts across inputs, such as linking an image of a street with the phrase ”urban road.” Multimodal alignment ensures that representations from different modalities map to a shared semantic space. Methods like contrastive learning align image and text pairs [81], while supervised fine-tuning improves cross-modal reasoning for downstream tasks [84]. In wireless contexts, aligning visual data with CSI or text with RF patterns is essential for applications like vision-RF fusion, semantic signal labeling, and instruction-following AI agents [85]. For example, time-aligned CSI and camera feeds enable accurate scene understanding and beam prediction. Cross-modal embedding spaces also support semantic compression, where audio is transcribed to text and transmitted at low bitrate requirement, improving spectral efficiency.

### v4

@@ 2.5 Fine-tuning
Fine-tuning pre-trained models with domain-specific data enhances performance for particular communication environments, which is essential for 6G and beyond as it eliminates the need for retraining from scratch. However, fine-tuning large AI models on edge nodes faces challenges beyond memory usage, such as high computational demands, energy efficiency, and network latency [yu2024snake]. The frequent exchange of gradients or parameters can also strain distributed learning frameworks. Additionally, techniques are necessary to adapt base models to high-level commands or specialized domains like telecommunications. To overcome these challenges, several promising techniques are introduced.
@@ 2.5.1 Parameter-Efficient Fine-Tuning
Low-rank adaptation (LoRA) [hu2021lora] adapts large pre-trained models by introducing low-rank updates to the weight matrices, freezing the original model parameters and training only the low-rank matrices, denoted as
@@
where $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ represents a pre-trained weight matrix which remains frozen during adaptation. The update $\Delta\mathbf{W}$ is constrained to be low-rank by decomposing it into the product of two smaller matrices, $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$, where the rank $r\ll\min(d,k)$. Only the parameters of $\mathbf{A}$ and $\mathbf{B}$ are trained, drastically reducing the number of trainable parameters compared to updating the full $\mathbf{W}_{0}$ or an unconstrained $\Delta\mathbf{W}$. This reduces the number of trainable parameters, making it ideal for resource-constrained environments. Variants like sparse LoRA [zhu2023sira] enhance efficiency using sparse expert modules, while weight-decomposed LoRA [liu2024dora] improves learning capacity by decomposing weights into magnitude and direction. LoRA also accelerates convergence in complex wireless datasets by simplifying the optimization landscape, preventing overfitting, and enabling faster adaptation to dynamic environments, improving model efficiency for tasks like traffic prediction and resource allocation.
@@
Prefix tuning [li2021prefixtuning] is a parameter-efficient method that adds a continuous, task-specific prefix sequence to the input or hidden layers of a model. This prefix sequence is not made up of real tokens but instead consists of learnable parameters. The prefix can influence the behavior of model without modifying the entire weights. This approach allows models to retain their pre-trained parameters while optimizing only the task-specific prefix, making it memory-efficient and computationally light.
@@
Prompt tuning [promptft] can be viewed as a simplified version of prefix tuning, where only a learnable prefix is added to the input text. This method has shown remarkable scaling capabilities; for sufficiently large models, prompt tuning alone can achieve performance comparable to full fine-tuning. By optimizing just the prompt parameters, prompt tuning provides an efficient way to adapt pre-trained models to specific tasks, minimizing the number of parameters that need to be updated. In wireless networks, prompt tuning can be used to efficiently optimize models for various real-time applications, reducing memory and computational costs, and ensuring practical deployment on edge devices.
@@ 2.5.3 Instruction Tuning
Instruction tuning [zhang2024instructiontuninglargelanguage] adapts a pre-trained model to better understand and follow natural language instructions across a wide range of tasks. Rather than training on task-specific data alone, the model learns from diverse instruction-response pairs, improving generalization and flexibility. In wireless networks, this allows operators to control or query the model using high-level commands such as “allocate bandwidth efficiently” or “detect congestion hotspots.” This approach reduces the need for hardcoded logic and enables more human-aligned behavior in communication environments. Instruction tuning is especially useful when tasks vary frequently or lack large labeled datasets.
@@ 2.6.1 Prompt-based Inference
LLMs can solve tasks without parameter updates by using prompts with instructions and examples, as in-context learning [brown2020language]. Instead of fine-tuning, the model generalizes from a few input-output demonstrations, enabling zero-shot and few-shot inference. This allows a single frozen model to adapt to various tasks by changing the prompt. Prompt engineering guides the model by designing text prompts, which can include system instructions, user queries, or domain-specific examples. This helps the model understand wireless contexts. Combined with instruction tuning [zhang2024instructiontuninglargelanguage], prompt engineering enables large models to support diverse use cases such as configuration, protocol analysis, and troubleshooting.
2.6.2 Chain of Thought
While basic prompts can handle simple queries, complex wireless problems benefit from multi-step reasoning. Chain of thought (CoT) prompting encourages the model to reason through intermediate steps before producing a final a answer [wei2022chain]. By appending phrases like reason step by step to the prompt, CoT enables more accurate and explainable results on tasks that require logical or numerical inference. This technique modifies the structure of few-shot exemplars. Instead of standard $(input,output)$ pairs, CoT exemplars take the form of (input,chain_of_thought,output)(input,chain\_of\_thought,output) triples, where the chain_of_thoughtchain\_of\_thought is a series of sentences outlining the reasoning process. Crucially, [wei2022chain] shows that CoT is an emergent ability of model scale; it provides significant performance gains only for models with approximately 100 billion or more parameters, and can even degrade performance on smaller models. This dependency on large-scale models makes CoT a powerful but resource-intensive tool, positioning it as a strategic choice for high-value, complex problems within wireless systems. In wireless systems, CoT is useful for anomaly detection, interference management, and resource optimization by guiding the model to iteratively explore the problem, effectively allocating more computation via token generation to tasks that require more reasoning steps.
@@ 2.6.3 Task Planning
Task planning is a more advanced reasoning paradigm that builds upon, but is distinct from CoT reasoning. While CoT focuses on generating a sequence of internal reasoning steps to solve a single, self-contained query, task planning operates as a higher-level, externally-focused action framework. Its primary purpose is to create and execute a structured workflow by decomposing a high-level goal into a series of explicit, executable sub-tasks.
The key distinction lies in the interaction with external tools and the environment. CoT is fundamentally an internal process, helping the model structure its thoughts to arrive at a final answer. In contrast, task planning acts as an orchestrator that not only plans the steps but also invokes external tools, application programming interfaces (APIs), or functions to perform actions at each stage. For example, to handle a high-level goal like ‘optimize the network for a major public event’, a task planning approach would create an executable plan, such as: first, calling a network monitoring API to get performance metrics; second, analyzing the retrieved data to confirm the bottleneck; and finally, calling a resource allocation API to reconfigure the network. Frameworks like HuggingGPT [shen2023hugginggpt] exemplify this by using an LLM as a controller to manage and invoke other specialized AI models to fulfill complex user requests. This pipeline-style, tool-augmented planning is essential for creating truly autonomous wireless systems that can sense, analyze, and act upon the network environment.
@@ 2.6.4 Retrieval Augmented Generation
Retrieval augmented generation (RAG) is an approach that enhances inference by coupling it with an external knowledge retriever [lewis2020retrieval]. In a RAG pipeline, a retriever first finds relevant documents or facts from a domain-specific datastore based on the query. The architecture features two main components: a neural retriever that finds relevant text passages for an input query and a generator that synthesizes an answer using the query and the retrieved content. This process treats the retrieved document $z$ as a latent variable that is marginalized to calculate the probability of generating a target sequence $y$ from an input $x$.
Two primary RAG models are commonly used, differing in how they utilize retrieved documents. The RAG-Sequence model uses the same retrieved document to generate the entire output sequence. The probability of generating sequence $y$ given input $x$ is calculated by summing the likelihoods over the top-$K$ retrieved documents:
$p_{\text{RAG-Sequence}}(y|x)\approx\sum_{z\in\text{top-k}(p_{\eta}(\cdot|x))}p_{\eta}(z|x)\prod_{i=1}^{N}p_{\theta}(y_{i}|x,z,y_{1:i-1}).$ (17)
On the other hand, the RAG-Token model can draw upon a different document for each token it generates, allowing for more dynamic synthesis of information. Its probability is calculated by marginalizing over the documents at each generation step:
$p_{\text{RAG-Token}}(y|x)=\prod_{i=1}^{N}\sum_{z\in\text{top-k}(p_{\eta}(\cdot|x))}p_{\eta}(z|x)p_{\theta}(y_{i}|x,z,y_{1:i-1}).$ (18)
In these formulations, $x$ is the input prompt, and $y=(y_{1},...,y_{N})$ is the generated output sequence of length $N$. The variable $z$ represents a retrieved document. The term $p_{\eta}(z|x)$ is the probability distribution of the retriever, which provides the likelihood of document $z$ being relevant to the input $x$. The summation $\sum_{z\in\text{top-k}(\cdot)}$ is performed over the set of top-$K$ documents returned by the retriever. Finally, $p_{\theta}(y_{i}|x,z,y_{1:i-1})$ is the probability of generating the next token $y_{i}$ given the original input $x$, the retrieved document $z$, and the previously generated tokens $y_{1:i-1}$, as modeled by the generator network with parameters $\theta$. These retrieved text snippets $z$ are provided as additional context to the generator, which integrates this evidence into its response. This mechanism effectively grounds the output in authoritative data, improving factual accuracy and reducing hallucinations [lewis2020retrieval].
@@
For example, TelecomRAG [telecomrag2023] applies this paradigm to telecommunication standards by building a knowledge base from 3GPP specification documents and enabling an LLM to answer protocol questions with precise, verifiable references. In wireless communications, RAG is useful for tasks like dynamic protocol lookup, fault diagnosis, and network troubleshooting [telerag]. An LLM agent can query a repository of network logs, configuration databases, or radiofrequency (RF) sensor readings and reason about network states in real-time. This yields grounded and trustworthy answers, especially important in rapidly evolving domains like wireless communications.
2.6.5 Mixture of Models
@@
Mixture of models uses multiple specialized models to tackle complex tasks by combining their expertise [moa]. In practice, queries are distributed to a set of models, each offering insights based on their specialized knowledge. The outputs are then evaluated and synthesized by a coordinating model, which selects or combines the most accurate responses. This approach helps improve accuracy, mitigate biases, and ensure more reliable decision-making in domain-specific applications, such as telecommunications [wang2025telemomconsensusdriventelecomintelligence], where precise expertise is critical.
2.6.6 Inference Optimization and Acceleration
Practical WLAM deployment is limited by inference efficiency, as serving is memory-bound due to the large and dynamically sized key-value (KV) cache. Conventional systems store the KV cache in contiguous memory, leading to severe fragmentation that can waste over 60% of the allocated memory, thereby limiting batch sizes and overall throughput. To solve this, vLLM [vllm] introduces PagedAttention, an algorithm that manages the KV cache in non-contiguous, fixed-size blocks, analogous to operation system paging. This design nearly eliminates memory fragmentation. The attention computation is adapted to operate on these non-contiguous blocks, where the output $o_{i}$ for a query $q_{i}$ is computed as follows:
$o_{i}=\sum_{j=1}^{\lceil i/B\rceil}\frac{\exp(q_{i}^{\top}K_{j}/\sqrt{d})}{\sum_{t=1}^{\lceil i/B\rceil}\exp(q_{i}^{\top}K_{t}\mathbf{1}/\sqrt{d})}V_{j}.$ (19)
Here $K_{j}$ and $V_{j}$ are blocks containing the key and value vectors, $B$ is the block size (a fixed number of tokens), and $d$ is the head dimension for scaling. By enabling much larger batch sizes, this method yields a 2 to 4 times throughput improvement over state-of-the-art systems. Frameworks like SGLang [sglang] further optimize complex, multi-call programs common in agentic workflows. For workloads with shared prefixes, its RadixAttention technique uses a radix tree to systematically reuse the KV cache across different requests, avoiding redundant computation. For structured generation (e.g., JSON), SGLang employs a compressed finite state machine that allows decoding of multiple deterministic tokens in a single forward pass. These co-designed runtime optimizations can increase throughput by up to 6.4 times on complex tasks.
@@ 2.7 Multimodal Models and Multimodal Alignment
The increasing complexity of wireless systems with their need to handle various types of data modalities (e.g., signals, images, text, and audio) has driven the development of multimodal models. These models integrate diverse modalities into a unified system, enabling more holistic reasoning and decision-making. By processing and aligning information across multiple types of input, multimodal models unlock powerful new applications in wireless communication, such as cross-modal inference, environment-aware communication, and sensor fusion. This section explores the key architectures, strategies, and challenges for enabling seamless integration of diverse data types.
2.7.1 Architectures and Fusion Strategies
Multimodal models integrate information from data types like text, images, and RF signals. The fundamental approach for processing multimodal data involves a two-stage process. First, modality-specific encoders are used to independently extract high-level features from each data type; for instance, a specific Transformer might process images while another specific Transformer handles language. Second, a fusion layer or mechanism combines these disparate feature streams into a unified, cross-modal representation. Architectures such as contrastive language-image pre-training [radford2021learning] and stable diffusion [stadiff] use modality-specific encoders and fusion layers, while models like Hunyuan-DiT [li2024hunyuandit] and Qwen-VL [qwenvl] apply these concepts to generative and question-answering tasks.
In wireless systems, the architectural choice reflects a trade-off between perception and efficiency. Multimodal Transformers leverage self-attention for complex perception tasks like sensing-assisted beam prediction [tian2023multimodaltransformerswirelesscommunications], while efficient models like vector-quantized variational autoencoders (VQVAE) are used for network-centric tasks such as CSI feedback optimization [Bocusvqvae].
As an example, the loss function for a multimodal VQVAE (MVQVAE) with $M$ modalities is designed to train encoders and decoders using a shared codebook, which is the key to fusing information. A representative loss function $L_{\text{MVQVAE}}$, adapted from the multimodal framework in [Bocusvqvae], is expressed as:
$L_{\text{MVQVAE}}=\sum_{m=1}^{M}\|\mathbf{x}_{m}-\hat{\mathbf{x}}_{m}\|^{2}_{2}+\sum_{m=1}^{M}\|\text{sg}(\mathbf{z}_{e,m})-\mathbf{z}_{q}\|^{2}_{2}+\beta\sum_{m=1}^{M}\|\mathbf{z}_{e,m}-\text{sg}(\mathbf{z}_{q})\|^{2}_{2}.$ (20)
The total loss sums three components over all $M$ modalities. The first term is the reconstruction loss between the original input $\mathbf{x}_{m}$ and the reconstruction $\hat{\mathbf{x}}_{m}$. The second term, the codebook loss, updates the shared codebook embedding $\mathbf{z}_{q}$ to match the encoder output $\mathbf{z}_{e,m}$. The third term, the commitment loss, is scaled by a hyperparameter $\beta$ and regularizes the encoder to produce vectors close to the codebook embeddings to minimize quantization error. The stop-gradient operator $\text{sg}(\cdot)$ isolates gradient updates to either the codebook or the encoders in the second and third terms, respectively, enabling end-to-end training.
Furthermore, these architectures implement various fusion strategies based on system requirements [shi2024radarcamerafusionobject]. Early fusion combines raw data at the input layer, ideal for cooperative sensing. Intermediate fusion merges features at middle layers, balancing performance and modularity. Late fusion combines final decisions from independent models, offering robustness for decentralized applications like federated learning.
2.7.2 Real-World Challenges and Alignment
The practical application of these methods in wireless environments faces significant hurdles from low-quality data [zhang2024multimodalfusionlowqualitydata]. Wireless signals are often noisy or incomplete due to sensor errors and interference. Data can also be imbalanced, causing models to over-rely on a dominant modality, while the quality of any single sensor may vary dynamically with environmental conditions. To overcome these issues, the fundamental goal is robust multimodal alignment. This process ensures that representations from different modalities map to a shared semantic space, for instance, through supervised fine-tuning [alayrac2022flamingo]. In wireless contexts, effectively aligning visual data with CSI or RF patterns is essential for applications like vision-RF fusion and semantic signal labeling, enabling more reliable and context-aware system performance [LLM4CP].
Table 3: Progress Analysis of WLAM with Its Design Principles.
Design Principle Current Progress & State-of-the-Art Existing Work Gaps & Future Directions
Adaptability to Dynamic Environments Use of RL, adaptive architectures (MoE, LNNs), and meta-learning for real-time adaptation to varying wireless conditions. Discussed applications in adaptive physical layer tasks (Sec 3.1), network resource allocation (Sec 3.2), and the Agentic AI-RAN paradigm for autonomous operation (Sec 3.4). Handling extreme non-stationary conditions (e.g., high mobility) remains a challenge. Future work needs more efficient on-device online training and zero-shot generalization to unseen scenarios.
Efficiency for Edge Deployment Advances in PEFT (e.g., LoRA), model compression, and efficient architectures (e.g., Mamba) with linear complexity. Survey covers distributed learning to reduce client load (Sec 4.1, 4.2), AirComp for communication efficiency (Sec 4.3), and next-gen sequence models (Sec 5.2.3). The performance-efficiency trade-off remains a key challenge. Future directions include better hardware-software co-design for edge AI and a stronger focus on energy efficiency, not just computational cost.
Multimodality and Data Diversity Development of large multimodal models for aligning diverse data (signals, text, images), including vision-aided wireless and RAG for knowledge fusion. Applications in ISAC for sensor fusion (Sec 3.1.3) and Semantic Communications for various data types (Sec 3.3). The need for multi-modal datasets is also identified (Sec 6.1.4). Challenges remain in the real-time fusion of unsynchronized data and ensuring robustness to missing modalities. There is a critical need for large-scale, standard wireless multi-modal datasets to advance future research.
Privacy and Security Use of federated learning for privacy, enhanced by techniques like differential privacy and lightweight physical layer security. Survey details the dual role of physical layer security for AI security (Sec 4.4) and AI for enhancing physical layer security (Sec 3.1.5). Core privacy methods like federated learning (Sec 4.2) and key threats (Sec 6.4) are discussed. Federated learning vulnerabilities (e.g., poisoning) and the high overhead of cryptography persist. Future directions include better balancing the privacy-security-performance trade-off and developing methods for auditing complex WLAM systems.